OpenStudy (anonymous):

triangle ABC~ triangle EDC by SSS. Name the correct proportion for the corresponding sides. A. 10/5=8/4=6/3 B. 10/4=8/3=6/5 C. 10/5=8/4=3/3 D. 10/5=4/8=6/3

OpenStudy (anonymous):

The answer is A all of those factors are divisible by two

OpenStudy (anonymous):

 @Luis_Rivera

OpenStudy (anonymous):

@e.cociuba

OpenStudy (anonymous):

A. 10/5=8/4=6/3